# Single-Image based unsupervised joint segmentation and denoising

### v1

@@ Abstract
In this work, we develop an unsupervised method for the joint segmentation and denoising of a single image. To this end, we combine the advantages of a variational segmentation method with the power of a self-supervised, single-image based deep learning approach. One major strength of our method lies in the fact, that in contrast to data-driven methods, where huge amounts of labeled samples are necessary, our model can segment an image into multiple meaningful regions without any training database. Further, we introduce a novel energy functional in which denoising and segmentation are coupled in a way that both tasks benefit from each other. The limitations of existing single-image based variational segmentation methods, which are not capable of dealing with high noise or generic texture, are tackled by this specific combination with self-supervised image denoising. We propose a unified optimisation strategy and show that, especially for very noisy images available in microscopy, our proposed joint approach outperforms its sequential counterpart as well as alternative methods focused purely on denoising or segmentation. Another comparison is conducted with a supervised deep learning approach designed for the same application, highlighting the good performance of our approach.
@@ 1 Introduction
Image denoising and segmentation are fundamental problems in image processing [37, 5, 23]. In many biomedical applications, such as fluorescence microscopy or transmission electron cryomicroscopy, one is interested in the segmentation of objects. However, training data for this task is typically scarce and hard to obtain due to the intrinsic complexity and high noise of such images as well as the long time required by experts to label them. Therefore, there is a need for unsupervised methods for tackling the two imaging tasks in a unified way. In this work, we propose such a framework, and apply it to a subset of a popular, public available dataset of microscopy images.
The objective of segmentation is to divide a given image into different, meaningful regions, while denoising describes the task of removing noise from a corrupted image. The main difficulty in noise removal is to flatten the unwanted, high frequency corruption, while preserving essential features such as edges. At first glance, denoising and segmentation are two different applications. Nevertheless, both tasks share relationships, as very similar models can be used to solve both problems [8]. As we demonstrate in this work, denoising and segmentation can benefit a lot from each other. By identifying edges segmentation guides the denoising process to preserve sharp structures while smoothing the unwanted high frequency residuals. Also, by removing unnecessary and misleading information from images, denoising helps and improves the segmentation accuracy.
@@
There exist at least two main kinds of approaches to tackle the two tasks individually. The first class of methods involves the minimisation of an energy functional within graph or variational frameworks. The second type of approaches that recently became popular considers deep learning techniques, especially convolutional neural networks [29]. In the following, we give a short overview of the most important and related variational and deep learning based methods.
1.1 Variational Methods
Standard imaging methods for segmentation and denoising are based on an energy functional that captures the desired characteristics of the output image. The energy functional typically consists of a data fitting term, and a regularisation term that encourages properties of the output image, such as smoothness or sparsity. The energy functional is then minimised using optimisation techniques such as gradient descent or proximal splitting algorithms.
Denoising
One of the best known variational model for image denoising is the Rudin-Osher-Fatemi (ROF) [36] model. This model improves the region-based Mumford-Shah  [31] functional that realises a piecewise smooth approximation of an input image. The ROF model and its extensions reduce noise by penalizing the total variation of the image. Such methods thus promote piecewise constant images with undesirable staircase effects in homogeneous regions and they are unable to recover image details and patterns with higher variation. In case of severe input noise, they provide poor denoising results as image contours are confused with noise [16].
On the other hand, the resulting image being piecewise constant, it can be used for segmentation, by choosing regions of the same value, or thresholding the image. More details about the link between the ROF based denoising models and segmentation can for example be found in [8].
Segmentation
In their seminal paper [11], Chan and Vese proposed to solve the MumfordShah problem with a levelset reformulation. Let us denote by $\Omega$ a bounded subset of $\mathbb{R}^{2}$ with Lipschitz boundary, where the given image $f:\Omega\rightarrow[0,1]$ is defined, and by $u\colon\Omega\rightarrow\{0,1\}$ the desired binary mask, separating $f$ into two different areas corresponding to the two mean intensity values $c_{1}$ and $c_{2}$, belonging to foreground and background region, respectively. In 2D, this framework involves a surface $\phi$ whose zero level represents the contour of interest, and the mask is obtained as $u(x)=H\left(\phi(x)\right),$ where $H(\cdot)$ is the Heaviside function. The proposed energy for binary segmentation is given by
[…]
$+\lambda\int_{\Omega}\lvert f(x)-c_{2}\rvert^{2}(1-H(\phi(x))dx,$
where $\lambda>0$ is a regularization parameter to tune. Slight modifications of this method have already been used in microscopy [34, 40], as it is well adapted to cell segmentation, where labeled data are scarce. which computes the intensity averages by using constant intensity information across the region. However, the Chan-Vese model computes the intensity averages by using constant information across the region, and thus is a global region-based model. Therefore, it does not deal well with intensity inhomogeneities and the presence of high noise. To mitigate this problem, many local region-based extensions of the active contour of the piecewise constant active contour model have been proposed [27, 45], but the methods remain sensible to the considered hand-crafted features and the initial contour. In another line of works, pre-filtering tools are considered to better prepare the image for segmentation [9, 28, 43] in a sequential pipeline: denoise before segment. In [7], a three-stage approach is proposed, consisting of smoothing, lifting, and segmentation using thresholding.
In our work, we tackle the afore-mentioned issues by introducing a generalized Chan-Vese segmentation functional including a robust data-fidelity term that is jointly learned with self supervised deep denoising techniques.
1.2 Deep Learning Methods
We now review some of the most relevant works using neural networks for the denoising and segmentation of images.
Denoising
While variational image denoising techniques focus on explicitly modeling data noise, modern deep learning approaches directly learn how to map the noisy image to its clean counterpart. In the literature, several types of deep learning based denoising methods can be found. In particular, supervised approaches require pairs of noisy images and corresponding clean ground truth data (see  [21, 44, 41]). However, the presence of such noisy and clean data pairs is rare in reality and often artificial, and it therefore makes such methods useless for practical applications.
To overcome the requirement of clean images, internal statistical methods (i.e. methods, where image patches from the same image are used for the noise reduction), have been introduced [46]. In [39], Ulyanov et al. exploit the fact that the internal structure of CNNs, inherently resonates with the distribution of natural images, and utilize this observation for image restoration without the need of additional training data. For each single image to restore, this method thus proposes to train a CNN to reconstruct the considered image. The idea is that an early stopping of the training allows to recover a regularised, denoised image. A different strategy is proposed in the Noise2Noise [25] method, where noisy image pairs are mapped to one another. The drawback of these type of methods is that it still relies on the availability of such pairs. In practice, even the acquisition of two noisy realisations of the same image content is often difficult [5]. To this end, self-supervised training methods operating on one single noisy image, such as Noise2Void [24], Noise2Self [3], and more recently Noise2Fast [26], are promising alternatives. This self-supervision is accomplished by excluding/masking the center (blind spot) of the receptive field of the network. In this type of training, it is assumed that the noise is pixelwise independent and that the true intensity of a pixel can be predicted from the local image context, with the exception of the blind spots mentioned previously [24]. In this work, We utilised Noise2Fast [26] in our study due to its optimal combination of computational speed and performance. The method itself will be explained in more detail in Sections 2, and 4.
Segmentation
Among existing deep learning based approaches addressing image segmentation, the U-Net [35] first introduced for microscopy cell segmentation, is one of the most successful network architectures. Next to it, we mention Mask-RCNN [20], a two-stage object detection and segmentation, extending the popular faster R-CNN architecture [17]. DeepLab [13] is a family of methods that use atrous convolution (also known as dilated convolution) to capture multi-scale context information. It has been shown to achieve state-of-the-art results on several segmentation benchmarks. Still, even the best existing methods offer plenty of scope for improvements, motivating further research in this field [21,23,7]. Their high performance comes along with a price to pay. A common trait to the mentioned approaches is their requirement for tremendous amounts of labeled ground truth training data, the creation of which is time-consuming and prone to subjective errors.
As already mentioned, for many applications such as microscopy, the available image data is very noisy and the available ground truth training data are scarce. It is thus of great interest to tackle both, the segmentation and the denoising, in a unified manner. In the following, we will review variational methods, as well as deep learning based approaches tackling both, the segmentation and the denoising in a joint manner.
1.3 Joint denoising and segmentation methods
In [6], Cai et al. design a model tackling the segmentation of images with a high level of noise or blurriness. To this end, they propose a variational approach, coupling an extension of the piecewise constant Mumford Shah model with an image restoration model, making it more robust in the processing of the given corrupted image $f$. In [14], the authors propose a variational approach for the joint reconstruction and segmentation. Therefore, they derive a model consisting of a total variation regularised reconstruction from undersampled data, and a Chan-Vese based segmentation. The authors show the improvement of joint reconstruction and segmentation performance compared to the sequential approach. In another work [33], Ramlau et. al illustrate that the Mumford–Shah level-et method can enhance the quality of reconstructed images and improve the accuracy of segmentation results.
In the context of microscopy data, purely deep learning based approaches dealing with both segmentation and denoising are [32] and [5]. In [32], Prakesh et al. demonstrated on various microscopy datasets that the use of self-supervised denoising priors improves the segmentation results, especially when only a few ground truth segmentation masks are available for training. In a similar work [5], the authors propose DenoiSeg, consisting of a U-Net for image segmentation, and the self-supervised denoising scheme Noise2Void [24], which are combined and trained with a common loss. The authors demonstrate, that the global optimisation outperforms the sequential counterpart, where the image is denoised first and then segmented. This method requires labeled data. To reach high denoising performance, a huge amount of noisy data is also required. Moreover, the loss function is just the sum of the segmentation and denoising losses. There exists no coupling between the two tasks in the objective to optimise. Segmentation can therefore benefit from the noise reduction, but the reverse is not possible.
To overcome these limitations, we a propose single-image method with a new joint loss that allows full interaction between segmentation and noise reduction.
1.4 Contributions
In this work, we propose a new model for joint image denoising and segmentation that combines advantages from variational models and deep learning. In contrast to the aforementioned state-of-the-art deep learning based methods which require a large cohort of labeled and clean training images, we obtain comparable results using only one single image. While supervised deep learning based methods are trained with hand-made annotations which are prone to subjective labelling errors, the proposed combination of a variational segmentation model with a self-supervised denoising CNN does not require any labeled data or a representative dataset leading to the elimination of pre-training.
We combine denoising and segmentation tasks in such a way, that both of them benefit from each other. This is a main difference between existing deep joint approaches such as [5], where the denoising task solely aims at improving segmentation. More specifically, we design two dedicated denoising networks for the foreground and background regions to improve the overall image denoising performance, and use the difference of the denoising performances in the two regions to find the segmentation mask.
Our method can be seen as a flexible generalization of existing Chan-Vese models. Standard Chan-Vese models as well as joint variational methods for segmentation and denoising  [6, 14] rely on the piecewise constant assumption of hand-crafted features. Thus, these methods are struggling with intensity inhomogeneties, textures or high noise levels. Further, methods of that kind strongly depend on the initialisation due to the non-convexity of the functional. In this work we propose to learn the structural patterns of different regions in the image without any prior information, by designing a novel energy functional, where feature information is captured in a natural way by a denoising network. The paper is organised as follows. We start with toy examples that illustrate the region specific denoisers as the motivation of our method in 1.5. We then formulate the two problems that we aim to solve and revise the necessary ingredients that make up our algorithm in Section 2. Our proposed new algorithm is described and analysed in Section 3, and the numerical implementation is presented in 4. Section 5 shows the application of the method to microscopy data. Further, we apply our proposed model to natural images, and demonstrate, that with manual guidance by the user roughly indicating the two regions, our method can successfully be applied to more complex problems. The paper ends with a conclusion and outlook to possible future work.
1.5 Motivation
@@
In the following, we give an intuition of how the two denoising and segmentation tasks we aim to solve are coupled in a way that both of them have a positive influence on each other. We present some toy examples, showing how segmentation can successfully guide the denoising process. In a toy example, we generate an image of size $\times 256$ that consists of stripe patterns that are aligned differently in different areas; see Figure 1. This image is further corrupted by manually adding Gaussian noise with a noise level of 50. Here, we use two linear neural networks (2 ”experts” respectively dedicated to the foreground and the background) consisting of a single convolutional layer with one filter of size $15\times 15$, which is trained using a slight modification of the Noise2Fast training strategy [26] described in Section 4. More precisely, we restrict the training of the network to the two regions of the image, respectively, by masking the loss function, and restricting the training on boxes of size $30\times 30$ which are depicted in Figure 1. We find that learned filters are adapted to the structure of the signal in the corresponding region. As a result, the error patterns have higher values in the region where the denoiser has not been trained. This provides the basis for exploiting region-specific denoising for segmentation. The experimental details for this toy example are provided in Section 5.
@@
The positive effect of segmentation on the denoising process is even more evident in the natural image shown in Figure 2. We used the network architecture proposed by the authors in [26], resulting in a non-linear neural network. First, the denoising network was trained and subsequently applied on the whole image. The second image shows the result obtained with two separately trained neural networks. This strategy yields a better visual result, which is further confirmed by PSNR values of 19.69 and 19.12, respectively. The noisy image has been generated by scaling the given clean RGB input image to $[0,1]$, and adding randomly distributed noise scaled with the maximum pixel value, ensuring that the noise is proportional to the image intensity. Here, we used a manually generated mask of the zebra and background region, and during training, computed the MSE restricted to the two different regions, respectively.
In the next section, we will fix the notation, formalize the problem, and describe the main ingredients that are used for our proposed method.
@@ 2 Problem Description
We now present in detail the background that is relevant for our algorithm. First we set our notations, and describe our proposed energy functional for our unified denoising and segmentation framework.
In the following, we denote by $\Omega\subset\mathbb{R}^{2}$ a bounded set with Lipschitz boundary, and by $\mathbb{F}$ a space of functions $f\colon\Omega\rightarrow\mathbb{R}^{d}$, with $d=1$ for grayscale images, and $d=3$ in the RGB case. We consider a given (noisy) image $f\in\mathbb{F}$, which we want to jointly denoise, and split up into $C$ different regions.
@@
The goal of image denoising is to recover a clean image $g$ from a noisy observation $f$ which follows an image degradation model $f=g+n$, where $n$ is the signal degrading noise which we want to remove.
Note that although other degradation types are possible, we assume an additive model here and specifically we will consider noise with an expected value of zero.
@@
Image segmentation refers to the process of automatically dividing an image into meaningful regions. Based on specific pre-defined characteristics of a given image $f\in\mathbb{F},$ one is interested in splitting the image domain into two (in the case of binary segmentation) regions $\Sigma,$ and $\Omega\setminus\Sigma$. In the case of multiclass segmentation, the objective is to build a partition $\Omega=\bigcup_{i=1}^{C}\Sigma_{i}$ of the image domain into $C$ disjoint regions (classes), where each of the regions $\Sigma_{1},\dots,\Sigma_{C-1}$ represents a specific structure of objects in $f$ and $\Omega\setminus(\Sigma_{1}\uplus\Sigma_{2}\uplus\dots\uplus\Sigma_{C-1})$ represents the background.
In this work, we address these two problems simultaneously by designing an energy functional in a way that both tasks benefit from each other. Next, we discuss the two main components from the literature that form the basis of our approach.

### v2

# Self2Seg: Single-Image Self-Supervised Joint Segmentation and Denoising

@@ Abstract
We develop Self2Seg, a self-supervised method for the joint segmentation and denoising of a single image. To this end, we combine the advantages of variational segmentation with self-supervised deep learning. One major benefit of our method lies in the fact, that in contrast to data-driven methods, where huge amounts of labeled samples are necessary, Self2Seg segments an image into meaningful regions without any training database. Moreover, we demonstrate that self-supervised denoising itself is significantly improved through the region-specific learning of Self2Seg. Therefore, we introduce a novel self-supervised energy functional in which denoising and segmentation are coupled in a way that both tasks benefit from each other. We propose a unified optimisation strategy and numerically show that for noisy microscopy images our proposed joint approach outperforms its sequential counterpart as well as alternative methods focused purely on denoising or segmentation.
@@ 1 Introduction
Image denoising and segmentation are fundamental problems in image processing [30, 5, 16]. Especially in many biomedical applications, such as fluorescence microscopy or transmission electron microscopy, one is interested in the segmentation of objects. Training data for this task is typically scarce and hard to obtain due to the intrinsic complexity and high noise of such images, as well as the long time required by experts to label them. Therefore, there is a need for unsupervised methods for tackling the two imaging tasks in a unified way. In this work, we propose such a framework by beneficially combining and extending ideas of the Chan-Verse segmentation model [9, 14] and self supervised learning [19, 17].
As we demonstrate in this paper joint self-supervised denoising and segmentation can benefit a lot from each other. By identifying individual regions, segmentation guides the denoising process to specific image structures. On the other hand, the adaptation of a denoiser to a specific region successfully guides the segmentation process.
@@
There exist at least two main kinds of approaches to tackle both tasks individually. The first class of methods involves the minimization of an energy functional within graph or variational frameworks. The second type of approaches that recently became popular considers deep learning techniques, especially those based on convolutional neural networks [22].
Before we start reviewing related works tackling the task of denoising and segmentation in a unified way, we provide the motivation behind the proposed strategy.
1.1 Motivation and Contributions
To overcome limitations of existing algorithms (see  1.2), in this paper we introduce a novel joint denoising and segmentation algorithm combining the strengths of variational models and deep learning.
@@ 1.1 Motivation and Contributions
We employ toy examples to demonstrate how segmentation effectively guides the denoising process. First, we generate a 256$\times$256 image featuring stripe patterns oriented differently across various regions, as depicted in Figure 1. Gaussian noise is manually added to corrupt the image. We utilize two linear neural networks, each serving as an ”expert” dedicated to either the foreground or background. These networks consist of a single convolutional layer with a $15\times 15$ filter, trained using a slightly modified version of the Noise2Fast strategy [19], detailed in Section 4. Specifically, we confine network training to the respective image regions by masking the loss function and restricting training to $30\times 30$ boxes, as illustrated in Figure 1. Not surprisingly, we observe that the learned filters adapt to the signal structure within their designated regions, leading to higher error patterns in untrained areas. This observation forms the basis for leveraging region-specific denoising for segmentation. Further experimental details for this toy example are available in Section 5.
@@
The positive effect of segmentation on the denoising process is even more evident in the natural image shown in Figure 2. We used the network architecture proposed by the authors in [19], resulting in a non-linear neural network. First, the denoising network was trained and subsequently applied to the whole image. The second image shows the result obtained with two separately trained neural networks. This strategy yields a better visual result, which is further confirmed by PSNR values of 19.69 and 19.12, respectively. The noisy image has been generated by scaling the given clean RGB input image to $[0,1]$, and adding randomly distributed gaussian noise scaled with the maximum pixel value, ensuring that the noise is proportional to the image intensity. Here, we used a manually generated mask of the zebra and background region, and during training, computed the mean squared error (MSE) restricted to the two different regions, respectively.
More specifically, we minimize the loss function
[…]
jointly over $u$, $\mathbf{D}_{F}$ and $\mathbf{D}_{B}$, taken from a specific class $\mathcal{U}\times\mathcal{D}_{F}\times\mathcal{D}_{B}$. Here $u$ and $1-u$ take values in $[0,1]$ and define the foreground and background regions and $\mathbf{D}_{F}$ and $\mathbf{D}_{B}$ are denoisers that should adapt to the two regions. Coupling a variational segmentation model with a self-supervised denoising CNN, overcomes the need for labeled data and pre-training. We integrate denoising and segmentation tasks to mutually enhance performance.
Specifically, we develop dedicated denoising networks for foreground and background regions, respectively. In that way, the overall denoising performance is improved by exploiting performance differences between the regions to derive the segmentation mask.
Conversely, we demonstrate that existing self-supervised denoising methods such as Noise2Self [1] or Noise2Fast [19] benefit from splitting up the image into regions by exploiting different structures and performing region-specific learning.
Unlike most previous deep learning methods for segmentation, our approach is self-supervised and achieves comparable results using just a single image. In contrast to existing approaches, we actually exploit the performance difference of denoisers which can vary across the different regions in the image, and show that by the way our approach is designed, both denoising and segmentation tasks have a positive influence on each other.
1.2 Related work
1.2.1 Joint denoising and segmentation
In order to improve noise robustness and handle intensity inhomogeneities in the Chan-Vese model, several local region-based extensions have been proposed in [20, 36, 27]. Nevertheless, these methods remain sensitive to manually crafted features and the initial contour. Pre-filtering tools to better prepare the image for segmentation are considered in [7, 21, 35]. The work by [24] proposes a segmentation-based image denoising algorithm for signal-dependent noise. After initial denoising, segmentation is applied to the pre-filtered image, and for each segment, the noise level is estimated. Subsequently, for each region, a separate denoiser is applied. In [6], a model tackling the segmentation of images with a high level of noise or blurriness is presented. To this end, they propose a variational approach, coupling an extension of the piecewise constant Mumford-Shah model [26] with an image restoration model, making it more robust in processing the given corrupted image. A variational approach for joint reconstruction and segmentation is proposed in [11], where a model consisting of a total variation regularized reconstruction from undersampled data and a Chan-Vese-based segmentation is used. The authors show the improvement of the joint reconstruction and segmentation performance compared to the sequential approach. In another paper by [29], the Mumford-Shah level set method is shown to improve the quality of reconstructed images and the accuracy of segmentation.
1.2.2 Self-supervised image denoising
In recent times, self-supervised deep learning techniques got increasing attention in the field of machine learning.
In [32], Ulyanov et al. exploit the fact that the internal structure of CNNs inherently resonates with the distribution of natural images, and utilize this observation for image restoration without the need for additional training data. For each single image to restore, this method thus proposes to train a CNN to reconstruct the considered image. The idea is that early stopping of the training allows for the recovery of a regularized, denoised image. A different strategy is proposed in Noise2Noise [18], where noisy image pairs are mapped to one another. The drawback of the latter method is that it still relies on the availability of such pairs. In practice, even acquiring two noisy realizations of the same image content is often difficult [5]. To this end, self-supervised training methods operating on one single noisy image, such as Noise2Void [17], Noise2Self [3], and more recently, Noise2Fast [19], have been proposed as promising alternatives. The self-supervision is accomplished by excluding/masking the center (blind spot) of the receptive field of the network. In this type of training, it is assumed that the noise is pixelwise independent and that the true intensity of a pixel can be predicted from the local image context, with the exception of the blind spots [17]. As one ingredient of our method we utilized the training strategy adapted from Noise2Fast [19], which is a variant of Noise2Self/Noise2Void. Noise2Fast offers the advantages of computational efficiency while maintaining strong performance. The method itself will be explained in more detail in Sections 2 and 4.
1.2.3 Learning methods for denoising and segmentation
In the context of microscopy data, purely deep learning-based approaches dealing with both segmentation and denoising are proposed in [28, 5]. In [28], it is demonstrated on various microscopy datasets that the use of self-supervised denoising priors improves the segmentation results, especially when only a few ground truth segmentation masks are available for training. In [5], the authors propose DenoiSeg, consisting of a U-Net for image segmentation and the self-supervised denoising scheme Noise2Void [17], which are combined and trained with a common loss. It is shown that global optimization outperforms the sequential counterpart, where the image is denoised first and then segmented. To reach high denoising performance, a huge amount of noisy data is required. Moreover, the loss function is just the sum of the segmentation and denoising losses. The fusion of the two tasks is achieved by sharing the network; the loss function does not incorporate any coupling between the two tasks. A slightly different approach is introduced in [23]. Nevertheless, this method requires ground truth data for both imaging tasks.
The paper is organized as follows: In Section 2, we formulate the problems and review algorithmic components. Section 3 details our proposed algorithm, followed by numerical implementation in Section 4. We apply the method to microscopy data in Section 5 and demonstrate its versatility on natural images with minimal user guidance. Finally, we conclude and discuss avenues for future research.
@@ 2 Problem Description
We will now fix the notation, formalize the problem, and describe the main ingredients that are used for the proposed unified denoising and segmentation method.
In the following, we denote by $\Omega\subset\mathbb{R}^{2}$ a bounded set with Lipschitz boundary, and by $\mathbb{F}$ a space of functions $f\colon\Omega\rightarrow\mathbb{R}^{d}$, with $d=1$ for grayscale images, and $d=3$ in the RGB case. We consider a given noisy image $f\in\mathbb{F}$, which we want to jointly denoise, and split up into $C$ different regions.
@@
The goal of image denoising is to recover a clean image $g$ from a noisy observation $f$ which follows an image degradation model $f=g+n$, where $n$ is the signal-degrading noise that we want to remove.
Note that although other degradation types are possible, we assume an additive model here, and specifically, we will consider random noise with an expected value of zero.
@@
Image segmentation refers to the process of automatically dividing an image into meaningful regions. Based on specific characteristics of a given image $f\in\mathbb{F}$, one is interested in splitting the image domain into two (in the case of binary segmentation) regions $\Sigma$ and $\Omega\setminus\Sigma$. In the case of multiclass segmentation, the objective is to build a partition $\Omega=\bigcup_{i=1}^{C}\Sigma_{i}$ of the image domain into $C$ disjoint regions (classes), where each of the regions $\Sigma_{1},\dots,\Sigma_{C-1}$ represents a specific structure of objects in $f$, and $\Omega\setminus(\Sigma_{1}\uplus\Sigma_{2}\uplus\dots\uplus\Sigma_{C-1})$ represents the background.
In this work, we address these two problems simultaneously by designing an energy functional in a way that both tasks benefit from each other. Next, we discuss the two main components that form the basis of our approach.